OpenStudy (drummmlers):

Add: 2a/a+b + a-b/a show steps please.

OpenStudy (anonymous):

first you need a common denominator therefore: \[(2a/a+b)*a =2a^2/(a^2+ab)\] \[((a-b)/a)*(a+b)=(a^2-b^2)/(a^2+ab)\] add them up you get: \[(3a^2-b^2)/a^2+ab)\]

OpenStudy (drummmlers):

thank you